OpenStudy (anonymous):

solve for x? (e^x+e^-x)/(e^x-e^-x)=3

OpenStudy (anonymous):

nice question!! e^2x/e^x + 1/e^x =3 (e^2x+1) / e^x = 3 e^2x + 1 = 3e^x e^2x - 3e^x + 1 = 0 e^x^2 -3e^x + 1 = 0 <---note this is a quadratic equation e^x = (3 +/- √(9-4))/2 <---so use the quadratic formula e^x = (3+√(5))/2 or (3-√(5))/2 x = ln ( (3+√(5))/2 ) or ln ( (3-√(5))/2 ) x = 0.96242365 or -0.96242365 hope this helps!! :)

OpenStudy (anonymous):

i guess it will be easier to apply componendo-dividendo

OpenStudy (lgbasallote):

@Rohangrr it seems you just covered the numerator..there's still a denominator

OpenStudy (anonymous):

i tried several times but i could not solve it.

OpenStudy (anonymous):

@Igbasallote ya!! but @MHAMMED u can solve for deno.. also !!

OpenStudy (anonymous):

Using the same way ? give me few second i am trying to understand what you did.

OpenStudy (lgbasallote):

oh wait..forgot to distribute...lemme repeat that...

OpenStudy (anonymous):

Guys really thankful for your help all of you

OpenStudy (lgbasallote):

rohan's first steps were right.. (e^2x/e^x + 1/e^x)/(e^2x/e^x - 1/e^x) = 3 (e^2x +1)/e^x/(e^2x - 1)/e^x = 3 (e^2x +1)/(e^2x - 1) = 3 e^2x + 1 = 3(e^2x - 1) e^2x + 1 = 3e^2x - 3 1+3 = 3e^2x - e^2x 4 = 2e^2x 2 = e^2x sqrt both sides... sqrt (2) = e^x ln both sides... x = ln (sqrt 2)

OpenStudy (anonymous):

well, i think my method is easier :|

OpenStudy (lgbasallote):

i think i got it wrong lol..wolfram agrees with @Arnab09

OpenStudy (anonymous):

all of you did well. thanks so much .

OpenStudy (experimentx):

cot hx = 3 rest is here http://en.wikipedia.org/wiki/Inverse_hyperbolic_function x = 1/2 (ln (4/2)) = 1/2 ln 2

OpenStudy (accessdenied):

@lgbasallote I don't think you were necessarily wrong, your answer just wasn't simplified. Since sqrt(2) = 2^(1/2) is in the logarithm, it can be taken outside using the power rule. So you'd still have ln(2)/2. :P

OpenStudy (lgbasallote):

oh oh yeah..glad i was right :DDD thanks @AccessDenied (Y) boost to the morale hehe